• Calculate the equilibrium concentration of all the species in a flask initially containing only 3.00 atm of NO2.

2NO2 (g) ⇌ N2O4 (g) Kp = 9.3 x 10-7
Answer #1

Kp = p(N2O4)/p(NO2)^2

9.3*10^-7 = x / (3.00-2x)^2

Since Kp is small, x will be small and can be ignored as compared to 3.00

So, the above expression becomes:

9.3*10^-7 = x / (3.00)^2

x = 8.4*10^-6 atm

So,

At equilibrium:

p(N2O4) = 8.4*10^-6 atm

p(NO2) = 3.00 atm

Answer:

p(N2O4) = 8.4*10^-6 atm

p(NO2) = 3.00 atm